OpenStudy (anonymous):

If f(x) = x/1 + x , what is the value of f(x + h)

OpenStudy (anonymous):

\[\frac{ x+h }{ 1-x-h}\]

ganeshie8 (ganeshie8):

\[f(\color{blue}{x}) = \dfrac{\color{blue}{x}}{1+\color{blue}{x}}\] simply replace \(\color{blue}{x}\) by \(\color{blue}{x+h}\) everywhere \[f(\color{blue}{x+h}) = ?\]

OpenStudy (anonymous):

\[\frac{ x+h }{ 1+x+h }\]

ganeshie8 (ganeshie8):

thats it!